OpenStudy (anonymous):

Can someone help me please??? I need this step by step! A radio station has a broadcast area in the shape of a circle with equation x2 + y2 = 5,625, where the constant represents square miles. a. Graph the equation and state the radius in miles. b. What is the area of the region in which the broadcast from the station can be picked up?

OpenStudy (amistre64):

the basic circle equation is: x^2 + y^2 = r^2 since r^2 = 5625, what is r?

OpenStudy (amistre64):

the area of a circle has the formula: A = pi r^2 and you already know r^2 ... so multiply it by pi